OpenStudy (anonymous):

If a job pays $24,000 the first year. During the next 14 years the salary increases by 5%. what is the total life salary over the 15 year period

OpenStudy (dumbcow):

\[S = 24000 (1 + 1.05+1.05^2 + ...1.05^{14})\] this is a geometric series with common ratio of 1.5 \[Sum = \frac{a_1(r^n -1)}{r-1}\] \[\rightarrow 24,000 \frac{1.05^{15} - 1}{1.05 - 1}\]

OpenStudy (anonymous):

thank you still working on a answer.

OpenStudy (dumbcow):

yw